OpenStudy (anonymous):

multiply, then simplify where possible. (3 multipled by sq. root of 7)^2 Please show step by step

myininaya (myininaya):

\[(3\sqrt{7})^2=3^2(\sqrt{7})^2=9(7)=63\]

jhonyy9 (jhonyy9):

9*7=63